• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name johna

    Log entry time 17:19:35 on May 2,2011

    Entry number 357839


    keyword=Max Juice values too low

    There were several trips in the first ~15 min while beam was being restored to both halls. I noticed that our Max Juice had been set from 127.5 back to 125 and that Hall C was at 175 (which was too low since they were planning to do tests at 175 and 180uA). Informed Dipangkar (Hall C RC) who called Jay Benesch. He called MCC and had them raised somewhat (127 for Hall A and 180 for Hall C), although they were not the source of the trips.

    So ours is not as high as it was yesterday, but it should be enough to avoid tripping when we have the half wave plate out (which gives us current spikes and Hall C current dips). Anyway, after ~10 min. tuning, the beam is better (independent of the Max Juice issue), and we'll have to wait a bit to see if it's stable enough.